# EEG-Driven 3D Object Reconstruction with Color Consistency and Diffusion Prior

### v1

@@ Abstract
EEG-based visual perception reconstruction has become a current research hotspot. Neuroscientific studies have shown that humans can perceive various types of visual information, such as color, shape, and texture, when observing objects. However, existing technical methods often face issues such as inconsistencies in texture, shape, and color between the visual stimulus images and the reconstructed images. In this paper, we propose a method for reconstructing 3D objects with color consistency based on EEG signals. The method adopts a two-stage strategy: in the first stage, we train an implicit neural EEG encoder with the capability of perceiving 3D objects, enabling it to capture regional semantic features; in the second stage, based on the latent EEG codes obtained in the first stage, we integrate a diffusion model, neural style loss, and NeRF to implicitly decode the 3D objects. Finally, through experimental validation, we demonstrate that our method can reconstruct 3D objects with color consistency using EEG.
@@ 1 Introduction
The human brain’s perception and interpretation of color is a complex process involving various neural mechanisms. Directly reconstructing and generating color-consistent 3D objects from brain activity data, such as EEG, represents a frontier exploration at the intersection of neuroscience and artificial intelligence. In recent years, text-based image generation techniques have made significant progress, greatly enhancing the quality and consistency of generated images, allowing precise representation of user-specified colors, shapes, and textures. Building on these advancements, the goal of this study is to explore the possibility of reconstructing color-consistent 3D objects using EEG signals, which represents an important step toward understanding and replicating the human visual perception process.
Figure 1: Our model reconstructs 3D objects using EEG signals.
@@
  • This paper presents a network model for EEG-to-3D color-consistent reconstruction, which, for the first time, attempts to extract color-consistent 3D perceptual features from EEG signals.
  • A two-stage neural encoding-decoding framework is proposed, enabling implicit neural encoding and decoding of color-consistent 3D objects driven by EEG signals.
  • Experimental results demonstrate that the proposed method can reconstruct color-consistent 3D objects using EEG signals.
@@ 2 Related Work
EEG-to-3D Generation
In recent years, significant progress has been made in text-to-3D object generation models  [22, 15, 16] , but EEG-to-3D object generation has not yet been fully explored. Inspired by text-to-3D methods such as  [15] and  [16], this paper attempts to utilize latent EEG codes to guide NeRF in reconstructing 3D objects through the shape priors of a diffusion model. To further investigate the brain’s ability to perceive color visual information from EEG, and drawing upon content and style transfer methods such as  [23] and  [24], we employ content and style losses to ensure that the reconstructed 3D objects maintain color consistency with ground truth (GT) images. Using this approach, we have, for the first time, achieved the reconstruction of color-consistent 3D objects from EEG signals, indirectly validating the neuroscientific theory that the human brain can perceive various types of visual information, such as color, shape, and texture, when observing objects.
@@ 3.2 Color Perception
The study in reference  [28] shows that the human brain efficiently encodes visual information through intricately designed pathways, where color and direction are structurally interrelated in brain signals. This indicates that these visual features are not processed independently but are systematically encoded together. According to reference  [29], using techniques such as intrinsic signal optical imaging, two-photon imaging, and electrophysiological recording, the study detailed the hue map structures of different visual brain areas, revealing the neural mechanisms underlying the formation of color perception space. It was found that as the processing hierarchy of the visual cortex increases, the neurons encoding color in the brain gradually coordinate, ultimately forming a balanced mechanism that matches our subjective perception of hues.
@@ 4.1 Overview
As shown in Figure 2, to reconstruct color-consistent 3D objects from EEG signals, we propose an EEG-to-3D architecture with color consistency and semantic awareness. The proposed method mainly consists of two components: 1) a semantic-aware implicit neural EEG encoder for 3D objects, and 2) a color-consistent implicit neural decoder for 3D objects. First, we utilize EEG signals for implicit neural encoding of 3D objects. In this stage, we obtain latent EEG features and class tokens through joint training tasks involving EEG signal reconstruction and classification, and these two components are merged and passed through a linear layer to generate EEG codes. Next, we use the EEG codes obtained in the previous stage to provide conditional decoding features for the reverse diffusion process of the diffusion model via a cross-attention mechanism, generating novel 2D views while using content and style losses to ensure that the colors of the novel 2D views remain as consistent as possible with GT. Finally, we use EEG codes to initialize the LDM to generate novel 2D views from different perspectives to optimize NeRF, enabling the implicit neural decoder to decode 3D objects.
4.2 Stage A: Multi-Task Joint Learning of EEG Encoder
During the initial phase, the process leverages EEG signals for the implicit neural encoding of 3D objects. Specifically, we first extract the regional semantic features of EEG by using a masked reconstruction model to capture the temporal features of EEG, thereby obtaining its regional information. Next, we perform a semantic classification task on EEG signals, enabling EEG encoder to learn semantic features. Through the EEG reconstruction task and semantic classification task, we capture the semantic features of regions, which serve as input for subsequent steps. Then, we explicitly introduce the semantic regional features of the original and reconstructed images by leveraging latent EEG codes containing latent information of semantic regions. We jointly fine-tune LDM to incorporate the performance of semantic regions, making the generated images more similar to GT in terms of regional and spatial relationships.
@@ 4.2.1 Reconstruction and Classification Tasks of EEG
EEG Classification Task
@@
4.2.2 Loss Function between EEG and Latent Diffusion Models
To endow LDM with regional semantic performance, we utilize the pre-trained SAM to obtain the regional semantic maps of visual stimulus images and their corresponding reconstructed images. The regional semantic loss is calculated by the cross-entropy loss function as follows,
${{\mathcal{L}}_{total}}={\lambda_{ldm}}{{\mathcal{L}}_{ldm}}+{\lambda_{region}}{{\mathcal{L}}_{region}},$ (3)
4.3 Stage B: EEG-to-3D with Color Consistency and Diffusion Priors
As Stage B, we utilize the reference view x generated by EEG and latent EEG codes to reconstruct NeRF, and constrain the novel view through a diffusion prior conditioned on the latent EEG codes. The proposed EEG-to-3D method is expected to simultaneously meet the following requirements: 1) It can generate 3D objects using EEG that closely resemble the rendering appearance of the reference view x; 2) Novel view renderings should exhibit semantic consistency with the reference view x, while also maintaining color consistency; 3) The generated 3D objects should possess well-defined geometric structures.
@@ 4.3 Stage B: EEG-to-3D with Color Consistency and Diffusion Priors
EEG-to-3D with Diffusion Priors
To ensure the semantic coherence of the EEG to 3D objects, additional constraints need to be applied to the novel view renderings. To address this challenge, we adopt a diffusion prior. Previous works on text-to-3D generation  [15, 16] have applied ${{\mathcal{L}}_{SDS}}$ to leverage text-conditioned diffusion models as 2D diffusion priors. In the method proposed in this paper, we employ a diffusion prior conditioned on latent EEG codes to optimize the views generated by NeRF, gradually refining them from blurry to sharp. In this case, to apply ${{\mathcal{L}}_{SDS}}$, we use the latent EEG codes obtained in the first stage as the latent prompt $y$, allowing us to perform ${{\mathcal{L}}_{SDS}}$ within the latent space of the diffusion model:
@@
EEG-Text Information Aligning with NeRF
According to ${{\mathcal{L}}_{SDS}}$ in DreamFusion, this function is used to measure the similarity between novel views and text prompts. However, in this work, 3D objects are generated based on EEG signals, thus requiring further alignment between the EEG signals and novel views from different perspectives. To achieve this, we leverage a pretrained CLIP text encoder to align the EEG signals with the novel views. The corresponding loss function is as follows:
${{\mathcal{L}}_{EEG-Text}}=-{{E}_{clip}}(T/I)\cdot\rho({{\varphi}_{latent}}(y))\cdot{{E}_{clip}}(g(\theta))$ (6)
Here, ${{E}_{clip}}(T/I)$ includes the CLIP image encoder and text encoder. Our objective is to align the text prompt of the novel view’s rendered image $g(\theta)$, the text prompt of the denoising image under the diffusion prior, and the EEG signals.
Prompt Visual Stimulus Images w/o ${{\mathcal{L}}_{color}}$ with ${{\mathcal{L}}_{color}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Prompt Visual Stimulus Images with ${{\mathcal{L}}_{color}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Latent EEG Codes [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Typical 3D objects generated by EEG-driven models based on the Things-EEG2 dataset  [21].
@@
${{\mathcal{L}}_{color}}=argmin\{{{\mathcal{L}}_{content}}(\varsigma(I),{{F}_{c}})+\lambda{{\mathcal{L}}_{style}}(\varsigma(I),{{F}_{s}})\}$ (7)
@@
Loss Function with Diffusion Priors and Color Consistency
The overall loss in the second stage can be represented by ${{\mathcal{L}}_{SDS}}$, ${{\mathcal{L}}_{EEG-Text}}$, and ${{\mathcal{L}}_{color}}$. In this work, the diffusion prior generated under the condition of latent EEG codes is used to guide NeRF in generating 3D objects with consistent color. ${{\mathcal{L}}_{SDS}}$ is used to optimize NeRF to achieve better geometric shapes and details, ${{\mathcal{L}}_{EEG-Text}}$ is employed to align the EEG, text, and image, and ${{\mathcal{L}}_{color}}$ is used to adjust the reconstructed color of the 3D objects to ensure they remain as consistent as possible with GT color, thereby enhancing the realism and visual accuracy of the 3D reconstruction.
@@ 5.1 Dataset and Implementation Details
In this work, the latent EEG codes obtained in Stage A are used to guide LDM in generating diffusion priors, with Stable Diffusion version 1.5 employed. To accelerate NeRF training and rendering in Stage B, the proposed method adopts Instant-NGP  [34]. We use ${{\mathcal{L}}_{SDS}}$, ${{\mathcal{L}}_{EEG-Text}}$, and ${{\mathcal{L}}_{color}}$ to compute the loss for generating 3D objects, and the Adam optimizer is used to update the model parameters, with the learning rate set to 0.001. Our hardware configuration utilizes an NVIDIA A100 GPU
@@
Full w/o ${{\mathcal{L}}_{color}}$ w/o ${{\mathcal{L}}_{EEG-Text}}$ w/o ${{L}_{region}}$
@@ 5.2 Results
5.2.1 Quality with 2D
As illustrated in Table 1, in Stage A, this study leverages EEG data combined with fine-tuning of LDM to reconstruct the generated 2D images. By accurately decoding the latent space that encapsulates semantic region information through EEG, the fine-tuned LDM successfully achieves a high-fidelity reconstruction of the original images.
5.2.2 Quality with 3D and Color Consistency
As shown in Table 2, this paper reconstructs 3D objects from EEG signals under two conditions: with and without style transfer loss. For the reconstruction of 3D objects with color consistency, the proposed method integrates latent EEG codes, semantically accurate reconstructed images, and style transfer to generate 3D objects with consistent color. These reconstructed 3D objects not only exhibit high geometric similarity to GT objects but also maintain consistency in color. This demonstrates that our model effectively captures the core features of 3D objects in the process of translating EEG signals into complex visual information, providing evidence that EEG signals can perceive and process various visual information such as color, shape, and texture.

### v3

# EEG-Driven 3D Object Reconstruction with Style Consistency and Diffusion Prior

@@ Abstract
Electroencephalography (EEG) -based visual perception reconstruction has become an important area of research. Neuroscientific studies indicate that humans can decode imagined 3D objects by perceiving or imagining various visual information, such as color, shape, and rotation. Existing EEG-based visual decoding methods typically focus only on the reconstruction of 2D visual stimulus images and face various challenges in generation quality, including inconsistencies in texture, shape, and color between the visual stimuli and the reconstructed images. This paper proposes an EEG-based 3D object reconstruction method with style consistency and diffusion priors. The method consists of an EEG-driven multi-task joint learning stage and an EEG-to-3D diffusion stage. The first stage uses a neural EEG encoder based on regional semantic learning, employing a multi-task joint learning scheme that includes a masked EEG signal recovery task and an EEG based visual classification task. The second stage introduces a latent diffusion model (LDM) fine-tuning strategy with style-conditioned constraints and a neural radiance field (NeRF) optimization strategy. This strategy explicitly embeds semantic- and location-aware latent EEG codes and combines them with visual stimulus maps to fine-tune the LDM. The fine-tuned LDM serves as a diffusion prior, which, combined with the style loss of visual stimuli, is used to optimize NeRF for generating 3D objects. Finally, through experimental validation, we demonstrate that this method can effectively use EEG data to reconstruct 3D objects with style consistency.
@@ 1 Introduction
A noninvasive brain-computer interface (BCI) system typically controls electronic devices through voluntary modulation of EEG signals. However, most current studies investigating the relationship between brain activity and 3D object visual imagery tasks primarily rely on functional magnetic resonance imaging (fMRI). For example, Mind-3D  [gao2023mind] successfully decoded 3D visual information from the brain using fMRI signals, demonstrating the feasibility of this challenging task. Therefore, existing fMRI studies provide theoretical foundations and research support for exploring different neuroimaging techniques and understanding the neural modulation in 3D object imagery tasks, especially in the context of EEG-based decoding of 3D visual information from the brain.
Figure 1: Our model reconstructs 3D objects using fMRI signals.
@@
  • We propose an EEG-based 3D object reconstruction framework with style-consistent semantic region awareness for reconstructing 3D objects that are consistent with the style of visual stimuli. The framework consists of an EEG multi-task joint learning stage and a style-semantic region-aware LDM fine-tuning and NeRF optimization stage. The former focuses on learning semantic- and location-aware latent EEG codes from EEG signals, while the latter uses these learned latent EEG codes as conditions to fine-tune LDM. The fine-tuned LDM serves as a diffusion prior, which, in combination with a style loss, optimizes NeRF and ultimately reconstructs style-consistent 3D objects.
  • We design a neural EEG encoder based on regional semantic learning, employing a multi-task joint learning scheme that includes a masked EEG signal recovery task and an EEG-based visual classification task. According to the visual attention shifting mechanism, the fixation regions of the human eye change over time. Thus, the recovery task helps the EEG encoder learn the spatial location information of objects by reconstructing missing EEG data, while the classification task aids in learning the semantic information of object regions.
  • This paper proposes a fine-tuning strategy for LDM with style-conditioned constraints and NeRF optimization. The strategy explicitly embeds semantic- and location-aware latent EEG codes and incorporates visual stimulus maps to fine-tune the LDM. The fine-tuned LDM serves as a diffusion prior, which, combined with the style loss of visual stimuli, is used to optimize NeRF. Finally, EEG data is employed to reconstruct 3D objects with color consistency.
@@ 2 Related Work
EEG-to-3D Object Generation
In recent years, significant progress has been made in text-to-3D object generation models  [22, 15, 16] , but EEG-to-3D object generation has not yet been fully explored. Inspired by text-to-3D methods such as  [15] and  [16], this paper attempts to utilize latent EEG codes to guide NeRF in reconstructing 3D objects through the shape priors of a diffusion model. To further investigate the brain’s ability to perceive color visual information from EEG, and drawing upon content and style transfer methods such as  [23] and  [24], we employ content and style losses to ensure that the reconstructed 3D objects maintain style consistency with ground truth (GT) images. Using this approach, we have, for the first time, achieved the reconstruction of style-consistent 3D objects from EEG signals, indirectly validating the neuroscientific theory that the human brain can perceive various types of visual information, such as color, shape, and texture, when observing objects.
@@ 3.2 Color Perception
According to reference  [29], using techniques such as intrinsic signal optical imaging, two-photon imaging, and EEG recording, the study detailed the hue map structures of different visual brain areas, revealing the neural mechanisms underlying the formation of color perception space. It was found that as the processing hierarchy of the visual cortex increases, the neurons encoding color in the brain gradually coordinate, ultimately forming a balanced mechanism that matches our subjective perception of hues.
@@ 4.1 Overview
As shown in Figure 2, to reconstruct style-consistent 3D objects from EEG signals, we propose an EEG-to-3D architecture with style consistency and semantic awareness. The proposed method mainly consists of two components: 1) a semantic-aware neural EEG encoder for 3D objects, and 2) a style-consistent neural decoder for 3D objects. First, we utilize EEG signals for neural encoding of 3D objects. In this stage, we obtain latent EEG features and class tokens through joint training tasks involving EEG signal reconstruction and classification, and these two components are merged and passed through a linear layer to generate latent EEG codes. Next, we use latent EEG codes obtained in the previous stage to provide conditional decoding features for the reverse diffusion process of the diffusion model via a cross-attention mechanism, generating novel 2D views while using content and style losses to ensure that the style of the novel 2D views remain as consistent as possible with GT. Finally, we use latent EEG codes to initialize LDM to generate novel 2D views from different perspectives to optimize NeRF, enabling the neural decoder to decode 3D objects.
4.2 Stage A: Multi-Task Joint Learning
During the initial phase, the process leverages EEG signals for neural encoding of 3D objects. Specifically, we first extract the regional semantic features of EEG by using a masked reconstruction model to capture the temporal features of EEG, thereby obtaining its regional information. Next, we perform a semantic classification task on EEG signals, enabling EEG encoder to learn semantic features. Through the EEG reconstruction task and semantic classification task, we capture the semantic features of regions, which serve as input for subsequent steps. Then, we explicitly introduce the semantic regional features of the original and reconstructed images by leveraging latent EEG codes containing latent information of semantic regions. We jointly fine-tune LDM to incorporate the performance of semantic regions, making the generated images more similar to GT in terms of regional and spatial relationships.
@@ 4.2.1 Reconstruction and Classification Tasks of EEG
EEG-based Visual Classification Task
@@
4.3 Stage B: Diffusion Prior and Style Consistency
As Stage B, we utilize the reference view x generated by EEG and latent EEG codes to reconstruct NeRF, and constrain the novel view through a diffusion prior conditioned on the latent EEG codes. The proposed EEG-to-3D method is expected to simultaneously meet the following requirements: 1) It can generate 3D objects using EEG that closely resemble the rendering appearance of the reference view x; 2) Novel view renderings should exhibit semantic consistency with the reference view x, while also maintaining style consistency; 3) The generated 3D objects should possess well-defined geometric structures.
@@ 4.3 Stage B: Diffusion Prior and Style Consistency
EEG-to-3D with Diffusion Prior
To ensure the semantic coherence of the EEG-to-3D objects, we adopt a diffusion prior to impose additional constraints on the novel view renderings. Previous works on text-to-3D generation  [15, 16] have applied ${{\mathcal{L}}_{SDS}}$ to leverage text-conditioned diffusion models as 2D diffusion priors. In the method proposed in this paper, we employ a diffusion prior conditioned on latent EEG codes to optimize the views generated by NeRF, gradually refining them from blurry to sharp. In this case, to apply ${{\mathcal{L}}_{SDS}}$, we use latent EEG codes obtained in the first stage as the latent prompt $y$, allowing us to perform ${{\mathcal{L}}_{SDS}}$ within the latent space of the diffusion model:
@@
EEG-Image Information Aligning with NeRF
According to ${{\mathcal{L}}_{SDS}}$ in DreamFusion, this function is used to measure the similarity between novel views and text prompts. However, in this work, 3D objects are generated based on EEG signals, thus requiring further alignment between the EEG signals and novel views from different perspectives. To achieve this, we leverage a pretrained CLIP image encoder to align the EEG signals with the novel views. The corresponding loss function is as follows:
${{\mathcal{L}}_{EEG\\ -CLIP}}=-{{E}_{CLIP}}(T/I)\cdot\rho({{\varphi}_{latent}}(y))\cdot{{E}_{CLIP}}(g(\theta))$ (3)
Here, ${{E}_{CLIP}}(T/I)$ includes the CLIP image encoder. By adding a similarity loss to the novel view image $g(\theta)$ obtained under the diffusion prior of EEG, our goal is to align the novel view image with the reference image.
Visual Stimulus Images w/o ${{\mathcal{L}}_{neuron-style}}$ with ${{\mathcal{L}}_{style}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Visual Stimulus Images with ${{\mathcal{L}}_{neuron-style}}$ 3D Views (${20}^{o}$) ${45}^{o}$ ${70}^{o}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: The method proposed in this study reconstructs 3D objects with typical style consistency using the Things-EEG2  [21] and fMRI-Image  [gao2023mind] datasets, respectively.
@@
${{\mathcal{L}}_{neuron-style}}=argmin\{{{\mathcal{L}}_{content}}(\varsigma(I),{{F}_{c}})+\lambda{{\mathcal{L}}_{style}}(\varsigma(I),{{F}_{s}})\}$ (4)
@@
Loss Function with Diffusion Prior and Style Consistency
The overall loss in the second stage can be represented by ${{\mathcal{L}}_{SDS}}$, ${{\mathcal{L}}_{EEG-CLIP}}$, and ${{\mathcal{L}}_{neuron-style}}$. In this work, the diffusion prior generated under the condition of latent EEG codes is used to guide NeRF in generating 3D objects with consistent style. ${{\mathcal{L}}_{SDS}}$ is used to optimize NeRF to achieve better geometric shapes and details, ${{\mathcal{L}}_{EEG-Image}}$ is employed to align the EEG, text, and image, and ${{\mathcal{L}}_{style}}$ is used to adjust the reconstructed style of the 3D objects to ensure they remain as consistent as possible with GT color, thereby enhancing the realism and visual accuracy of the 3D reconstruction.
Loss Function between EEG and Latent Diffusion Models
To endow LDM with regional semantic performance, we utilize the pre-trained Segment Anything (SAM)  [kirillov2023segment] to obtain the regional semantic maps of visual stimulus images and their corresponding reconstructed images. The regional semantic loss is calculated by the cross-entropy loss function as follows,
${{\mathcal{L}}_{ldm-region}}={\lambda_{ldm}}{{\mathcal{L}}_{ldm}}+{\lambda_{region}}{{\mathcal{L}}_{region}},$ (7)
@@ 5.1 Dataset and Implementation Details
fMRI-Image Dataset
The fMRI dataset comes from Mind-3D  [gao2023mind], and in the supplementary materials, the methods proposed in this paper are compared with Mind-3D. The methods in this paper reconstruct 3D objects with consistent style.
@@
In this work, the latent EEG codes obtained in Stage A are used to guide LDM in generating diffusion priors, with Stable Diffusion version 1.5 employed. To accelerate NeRF training and rendering in Stage B, the proposed method adopts Instant-NGP  [34]. We use ${{\mathcal{L}}_{SDS}}$, ${{\mathcal{L}}_{EEG-Image}}$, and ${{\mathcal{L}}_{neuron-style}}$ to compute the loss for generating 3D objects, and the Adam optimizer is used to update the model parameters, with the learning rate set to 0.001. Our hardware configuration utilizes an NVIDIA A100 GPU
@@
Full w/o ${{\mathcal{L}}_{neuron-style}}$ w/o ${{\mathcal{L}}_{EEG-CLIP}}$ w/o ${{L}_{region}}$
@@ 5.2 Results
5.2.1 Quality with 3D and Style Consistency
As shown in Table 1 and  2, this paper reconstructs 3D objects from EEG signals under two conditions: with and without style transfer loss. For the reconstruction of 3D objects with style consistency, the proposed method integrates latent EEG codes, semantically accurate reconstructed images, and style transfer to generate 3D objects with consistent style. These reconstructed 3D objects not only exhibit high geometric similarity to GT objects but also maintain consistency in style. This demonstrates that our model effectively captures the core features of 3D objects in the process of translating EEG signals into complex visual information, providing evidence that EEG signals can perceive and process various visual information such as color, shape, and texture.